OpenStudy (anonymous):

two complementary angles have a difference of 18 degrees. find their measures.

OpenStudy (anonymous):

Let the Complementary angles be x and y X+Y = 90 X - Y = 18 2x = 108 x = 54 y = 90 -x = 90 - 54 = 36